OpenStudy (anonymous):

How many liters of SO3 will be produced when 25 L of sulfur dioxide reacts with 75 L of oxygen? All gases are at STP.

OpenStudy (anonymous):

i know 2SO2 + O2 --> 2S03

OpenStudy (abmon98):

\[2SO2 + O2\rightarrow2SO3\] STP is short for Standard Temperature and Pressure, which is defined to be 273 K (0° Celsius) and 1 atm pressure. STP describes Standard Conditions. STP is often used for measuring gas density and volume using the Ideal Gas Law. Here, 1 mole of ideal gas occupies 22.4 L Number of moles*22.4=Volume of Gas Number of moles of O2=75/22.4=3.34 Number of moles of SO2=25/22.4=1.11 Once reactant is used up before another runs out. This reactant is known as the limiting reactant. We use mole ratio according to the given balanced chemical equation. 2 moles of SO2 react with 1 mole of O2 to produce 2 moles of SO3. 2:1:2 Before we use mole ratio you need to know that 3.34 and 1.11 moles are the actual amounts present by each reactant. using mole ratio to find the limiting reagent is just a theoretical proof and a comparison between the theoretical and the actual moles. present. Use mole ratio based on the O2 moles present. 2:1 x:3.34 x represents the "theoretical" number of moles of SO2 required to react with 3.34 moles of O2 Do a Cross multiplication to find out x. x=3.34*2/1=6.68 moles of SO2 and the actual number of moles of SO2 present is 1.11, this shows that SO2 is in inadequate amount and therefore its the limiting reagent. Now use the mole ratio again between SO2 and SO3, its 1:1 both share the same number of moles. so 1.11 is the number of moles of SO3 produced by 1.11 moles of SO2 reacting. Again, use the equation Number of moles*22.4=Volume (L) 1.11*22.4=24.864 L

OpenStudy (somy):

@Abmon98 lol all fully done again xD

OpenStudy (abmon98):

@Somy hahaha, to be honest i did that on purpose, i stopped logging in to OS regularly as before so i replied with a detailed which i think is correct and understandable?

OpenStudy (somy):

i think it is :33 @Abmon98